STAR is a parallelogram, if ST is 2x + 5 and RA is 16. What is the value of x?
Answer by ikleyn(52788)

.

ST and AR are opposite sides of the parallelogram STAR.


They are congruent and have equal lengths.  So

    ST = AR,  or  2x + 5 = 16.


It implies  

    2x = 16 - 5 = 11.

     x          = 11/2 = 6.5.    ANSWER

Solved.